OpenStudy (anonymous):

evaluate without a calculator: cot(tan^-1 (1/2))

OpenStudy (anonymous):

first evaluate tan^-1(1/2) to evaluate an inverse trig function like this you ask yourself, the tangent of what is equal to 1/2?

OpenStudy (anonymous):

so tan(x) = 1/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do i find that?

OpenStudy (shubhamsrg):

 if you notice, you see you'll have to find the value of cosx

OpenStudy (anonymous):

ok, so sin(x) = sqrt of 3/1

OpenStudy (shubhamsrg):

nopes..recheck..

OpenStudy (anonymous):

cos (x) = sqrt 3 / 2

OpenStudy (anonymous):

sqrt 5

OpenStudy (shubhamsrg):

yes,,it'll be sqrt 5.. now cosx = 2/sqrt5 and hence your ans.

OpenStudy (anonymous):

of the whole thing?

OpenStudy (kainui):

Remember, cotangent is Adjacent/Opposite while tangent is Opposite/Adjacent. By doing one and then the other, you're really just flip-flopping the fraction! No need to get crazy here with the math.

OpenStudy (anonymous):

oh yea... dur check this out \[ \tan^{-1} (\sqrt{3}) = \pi / 3\] \[\cot(\pi / 3) = \frac{ 1 }{ \sqrt{3} }\] just flipped it... thanks Kainui!

OpenStudy (kainui):

Another way to look at it is that cot(x)=1/tan(x) so you can really just say this problem is doing: \[\frac{ 1 }{ \tan(\tan^{-1}(x)) }\]

OpenStudy (kainui):

So yeah, you don't really have to plug this one into a calculator lol.